OpenStudy (anonymous):

Find the solution to each system by the substitution method. Check your answers for exercises 15-18. 3x-5y=7 x-4y=-7

OpenStudy (anonymous):

So if you add 4y to second would be x=-7+4y now plug this into first equation So 3(-7+4y)-5y=7 once u get y plug it in to get x

OpenStudy (anonymous):

What then do u get ?

OpenStudy (anonymous):

For my y I got this if you can tell me if it was right y=28/7=4. So y=4. What is the next formula for x can you please help me?

OpenStudy (anonymous):

I'm so sorry I just having problems I get don't understand it. sorry again

OpenStudy (anonymous):

ok if I did that and I got for x is 43 is that right. Thank you for helping me out on this.

OpenStudy (anonymous):

Ok here equation is x-4(4)=-7

OpenStudy (anonymous):

X-16=-7 add 16 to -7 to isolate x and answer should be x=9

OpenStudy (anonymous):

ok thank you so much for your help.

OpenStudy (anonymous):

:) no problem

OpenStudy (anonymous):

Now to check your answer plug in the x and y values So 3(9)-5(4)=7 ?